OpenStudy (anonymous):

evaluate the expression described belowif the variable givenis 9. two times the difference of one-third of a number and three

OpenStudy (anonymous):

please help me

OpenStudy (pradipgr817):

evaluate the expression described belowif ....?

OpenStudy (anonymous):

\[2(1/3n-3)=2(1/3*9-3)=0\]